英語 (English)

発信型英語 (Communicative English) (総科 ((社創 C)2 年))

石田メグ · 非常勤講師 2 単位前期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This English course is more about using the knowledge of English in conversation and written form. In this way, students will be ever ready for service in a global community. Japanese students will be prepared to face a global field without hesitation or insecurity. This course focuses on conversational English, maximizing speaking time spent in the classroom.

【授業の概要】 Students will create their own conversations about topics familiar to them. Classes will include whole, small group and pair activities. Students will be required to take risks and assess their own progress as the course develops. Students who take risks will do well in this class.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students who complete this course should feel more confident when

speaking or writing in English. 【授業の計画】

1. OUTLINE

2. I think I’d like that job

3. What’s playing

4. What are you going to do?

5. How much is this?

6. How was it?

7. How do you make it? 【教科書】 No text is needed. Handouts will be prepared. Please bring your English

Japanese Dictionary. 【参考書等】 [参考資料] 【成績評価の方法】 Attendance 10% Class participation 40% Student self-evaluation

10% Final Interview 40% 【再試験の有無】有【受講者へのメッセージ】 Come to class and participate and you will succeed! The contract

for this class follows University guidelines. If more than 4 classes are missed there will be no pass grade for this course. Students more than 20 minutes late are considered absent. If students are late twice it equals one absence. Students should bring a notebook, black and red pens, a pencil, the textbook, and a Japanese-English dictionary with them to class. Do not use ketai phones in class. Be present and participate. The contents of this class may vary due to the interests and also the level of the students.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220888 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 石田 .

English

Communicative English (総科 ((社創 C)2 年)) Meg Ishida · Part-time Lecturer 2 units 前期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This English course is more about using the knowledge of English in conversation and written form. In this way, students will be ever ready for service in a global community. Japanese students will be prepared to face a global field without hesitation or insecurity. This course focuses on conversational English, maximizing speaking time spent in the classroom.

Outline⟩ Students will create their own conversations about topics familiar to them. Classes will include whole, small group and pair activities. Students will be required to take risks and assess their own progress as the course develops. Students who take risks will do well in this class.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students who complete this course should feel more confident when speaking

or writing in English.

Schedule⟩

1. OUTLINE

2. I think I’d like that job

3. What’s playing

4. What are you going to do?

5. How much is this?

6. How was it?

7. How do you make it? Textbook⟩ No text is needed. Handouts will be prepared. Please bring your

English Japanese Dictionary. Reference⟩ [参考資料] Evaluation Criteria⟩ Attendance 10% Class participation 40% Student self-

evaluation 10% Final Interview 40% Re-evaluation⟩ 有 Message⟩ Come to class and participate and you will succeed! The contract for

this class follows University guidelines. If more than 4 classes are missed there will be no pass grade for this course. Students more than 20 minutes late are considered absent. If students are late twice it equals one absence. Students should bring a notebook, black and red pens, a pencil, the textbook, and a Japanese-English dictionary with them to class. Do not use ketai phones in class. Be present and participate. The contents of this class may vary due to the interests and also the level of the students.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220888 Contact (Office-Hour, Room, E-mail)⟩

⇒ Ishida .